# On the discontinuities of Hausdorff dimension in generic dynamical Lagrange spectrum

### v1

@@ Abstract.
Let $\varphi_{0}$ be a $C^{2}$-conservative diffeomorphism of a compact surface $S$ and let $\Lambda_{0}$ be a mixing horseshoe of $\varphi_{0}$. Given a smooth real function $f$ defined in $S$ and some diffeomorphism $\varphi$, close to $\varphi_{0}$, let $\mathcal{L}_{\varphi,f}$ be the Lagrange spectrum associated to the hyperbolic continuation $\Lambda(\varphi)$ of the horseshoe $\Lambda_{0}$ and $f$. We show that, for generic choices of $\varphi$ and $f$, if $L_{\varphi,f}$ is the map that gives the Hausdorff dimension of the set $\mathcal{L}_{\varphi,f}\cap(-\infty,t)$ for $t\in\mathbb{R}$, then there are at most two points that can be limit of a infinite sequence of discontinuities of $L_{\varphi,f}$.
@@ 1.1. Classical spectra
:=$\mathcal{L}:=\{\ell(\alpha):\alpha\in\mathbb{R}-\mathbb{Q}\ \text{and}\ \ell(\alpha)<\infty\}$
@@
Similarly, given a real quadratic form $q(x,y)=ax^{2}+bxy+cy^{2}$, let $\Delta(q)=b^{2}-4ac$ its discriminant. We define the Markov spectrum as follows
:=$\mathcal{M}:=\left\{\frac{\sqrt{\Delta(q)}}{\inf\limits_{(x,y)\in\mathbb{Z}^{2}-\{(0,0)\}}|q(x,y)|}<\infty:q\ \text{is indefinite and}\ \Delta(q)>0\right\}.$
@@
  • Hall showed that $\mathcal{L}$ (and then $\mathcal{M}$) contain a half-line and Freiman determined the biggest half-line contained in the spectra, namely $[c,+\infty)$ where
@@
For our purposes, it is worth to point out here that the Lagrange and Markov spectra have the following dynamical interpretation in terms of the continued fraction algorithm: Denote by $[a_{0},a_{1},\dots]$ the continued fraction $a_{0}+\frac{1}{a_{1}+\frac{1}{\ddots}}$. Let $\Sigma=\mathbb{N}^{\mathbb{Z}}$ the space of bi-infinite sequences of positive integers, $\sigma:\Sigma\to\Sigma$ be the left-shift map $\sigma((a_{n})_{n\in\mathbb{Z}})=(a_{n+1})_{n\in\mathbb{Z}}$, and let $f:\Sigma\to\mathbb{R}$ be the function
@@
=$\mathcal{L}=\left\{\limsup_{n\to\infty}f(\sigma^{n}(\underline{\theta}))<\infty:\underline{\theta}\in\Sigma\right\}\quad\textrm{and}\quad\mathcal{M}=\left\{\sup_{n\to\infty}f(\sigma^{n}(\underline{\theta}))<\infty:\underline{\theta}\in\Sigma\right\}.$
In the sequel, we consider the natural generalization of this dynamical version of the classical Lagrange and Markov spectra in the context of horseshoes111i.e., a non-empty compact invariant hyperbolic set of saddle type which is transitive, locally maximal, and not reduced to a periodic orbit (cf. [16] for more details). of smooth diffeomorphisms of compact surfaces.
@@ 1.2. Dynamical spectra
Let $\varphi:S\rightarrow S$ be a diffeomorphism of a $C^{\infty}$ compact surface $S$ with a mixing horseshoe $\Lambda$ and let $f:S\rightarrow\mathbb{R}$ be a differentiable function. Following the above characterization of the classical spectra, we define the maps $\ell_{\varphi,f}:\Lambda\rightarrow\mathbb{R}$ and $m_{\varphi,f}:\Lambda\rightarrow\mathbb{R}$ given by $\ell_{\varphi,f}(x)=\limsup\limits_{n\to\infty}f(\varphi^{n}(x))$ and $m_{\varphi,f}(x)=\sup\limits_{n\in\mathbb{Z}}f(\varphi^{n}(x))$ for $x\in\Lambda$ and call $\ell_{\varphi,f}(x)$ the Lagrange value of $x$ associated to $f$ and $\varphi$ and also $m_{\varphi,f}(x)$ the Markov value of $x$ associated to $f$ and $\varphi$. The sets222we omit the reference to the horseshoe $\Lambda$ because in our context it will be always determined by the diffeomorphism $\varphi.$
@@
are called Lagrange Spectrum of $(\varphi,f)$ and Markov Spectrum of $(\varphi,f)$.
@@
In order to do that, we will explore the combinatorial nature of $\varphi|_{\Lambda}$ and its connection with the unstable and stable Cantor sets associated to $\Lambda$. More specifically, fix a Markov partition $\{R_{a}\}_{a\in\mathcal{A}}$ of $\Lambda$ with sufficiently small diameter consisting of rectangles $R_{a}\sim I_{a}^{u}\times I_{a}^{s}$ delimited by compact pieces $I_{a}^{s}$, $I_{a}^{u}$, of stable and unstable manifolds of certain points of $\Lambda$, see [16] theorem 2, page 172. The set $\mathcal{B}\subset\mathcal{A}^{2}$ of admissible transitions consist of pairs $(a,b)$ such that $\varphi(R_{a})\cap R_{b}\neq\emptyset$; so, we can define the transition matrix $B$ by
@@
Let $\Sigma_{\mathcal{A}}=\left\{\underline{a}=(a_{n})_{n\in\mathbb{Z}}:a_{n}\in\mathcal{A}\ \text{for all}\ n\in\mathbb{Z}\right\}$ and consider the homeomorphism of $\Sigma_{\mathcal{A}}$, the shift, $\sigma:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{A}}$ defined by $\sigma(\underline{a})_{n}=a_{n+1}$. Let $\Sigma_{\mathcal{B}}=\left\{\underline{a}\in\Sigma_{\mathcal{A}}:b_{a_{n}a_{n+1}}=1\right\}$, this set is closed and $\sigma$-invariant subspace of $\Sigma_{\mathcal{A}}$. Still denote by $\sigma$ the restriction of $\sigma$ to $\Sigma_{\mathcal{B}}$, the pair $(\Sigma_{\mathcal{B}},\sigma)$ is a subshift of finite type, see [6] chapter 10. The dynamics of $\varphi$ on $\Lambda$ is topologically conjugate to the sub-shift $\Sigma_{\mathcal{B}}$, namely, there is a homeomorphism $\Pi:\Lambda\to\Sigma_{\mathcal{B}}$ such that $\varphi\circ\Pi=\Pi\circ\sigma$.
As we generally will deal with sequences, we transfer the function $f$ from $\Lambda$ to a function (still denoted $f$) on $\Sigma_{\mathcal{B}}$. In this way, we set
@@
Recall that the stable and unstable manifolds of $\Lambda$ can be extended to locally invariant $C^{1+\alpha}$ foliations in a neighborhood of $\Lambda$ for some $\alpha>0$. Using these foliations it is possible define projections $\pi^{u}_{a}:R_{a}\rightarrow I^{s}_{a}\times\{i^{u}_{a}\}$ and $\pi^{s}_{a}:R_{a}\rightarrow\{i^{s}_{a}\}\times I^{u}_{a}$ of the rectangles into the connected components $I^{s}_{a}\times\{i^{u}_{a}\}$ and $\{i^{s}_{a}\}\times I^{u}_{a}$ of the stable and unstable boundaries of $R_{a}$, where $i^{u}_{a}\in\partial I^{u}_{a}$ and $i^{s}_{a}\in\partial I^{s}_{a}$ are fixed arbitrarily. In this way, we have the unstable and stable Cantor sets
@@
Moreira’s theorem of [11] was generalized first in [1] in the context of conservative diffeomorphism with some horseshoe with Hausdorff dimension smaller than $1$ and later was removed the condition on the dimension of the horseshoe in [9]. More specifically, the authors proved that for typical choices of the dynamic and of the real function, the intersections of the corresponding dynamical Markov and Lagrange spectra with half-lines $(-\infty,t)$ have the same Hausdorff dimension, and this defines a continuous function of $t$ whose image is $[0,\min\{1,D\}]$, where $D$ is the Hausdorff dimension of the horseshoe.
Our main theorem (cf. Theorem 1.1 below) is quite related to the result of the previous paragraph but, in our case, we will work away from the two points that determine “the canonical interval” where $L_{\varphi,f}$ can have a discontinuity. Here, we drop the hypothesis of the neighborhood of the initial conservative diffeomorphism be in the space of conservative diffeomorphisms. However, we can only conclude finiteness of the number of discontinuities but not continuity else.
@@ 1.3. Statement of the main theorem
  • If $HD(\Lambda)<1$ then $L_{\varphi,f}$ has finitely many discontinuities in any closed sub interval $I\subset I_{\varphi,f}$ that doesn’t contain $c_{\varphi,f}.$
  • If $HD(\Lambda)\geq 1$ then $L_{\varphi,f}$ has finitely many discontinuities in any closed sub interval $I\subset I_{\varphi,f}$ that doesn’t contain neither $c_{\varphi,f}$ nor $\tilde{c}_{\varphi,f}.$
@@
If $HD(\Lambda_{0})<1$, then by choosing $\mathcal{U}$ small, given $\varphi\in\mathcal{U}^{*}$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and $\epsilon>0$ the function $L_{\varphi,f}$ has finitely many discontinuities in the interval $[c_{\varphi,f}+\epsilon,\infty)$. Therefore, $c_{\varphi,f}$ is the only possible limit of an infinite sequence of discontinuities of $L_{\varphi,f}$.
@@
If $HD(\Lambda_{0})>1$, then by choosing $\mathcal{U}$ small, given $\varphi\in\mathcal{U}^{*}$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and $\epsilon>0$ small, the function $L_{\varphi,f}$ has finitely many discontinuities in the interval $[c_{\varphi,f}+\epsilon,\tilde{c}_{\varphi,f}-\epsilon].$ Therefore, $c_{\varphi,f}$ and $\tilde{c}_{\varphi,f}$ are the only possible limits of an infinite sequence of discontinuities of $L_{\varphi,f}$.
@@ 2.1. Stable and unstable dimensions
Note that if $x\in R(\theta;s)\cap\Lambda$ then the symbolic representation of $x$ is in the way $\Pi(x)=(\dots,a_{s_{1}}\dots a_{s-1};a_{s},a_{s+1}\dots a_{s_{2}}\dots)$, where the letter following to ; is in the 00 position of the sequence.
In our context of dynamically defined Cantor sets, we can relate the length of the unstable and stable intervals determined by an admissible word to its length as a word in the alphabet $\mathcal{A}$ via the bounded distortion property that let us conclude that for some constant $c_{1}>0$
@@
We write $r^{(u)}(\alpha)$ for the unstable scale of $\alpha$, that is, $r^{(u)}(\alpha)=\lfloor\log(1/\lvert I^{u}(\alpha)\rvert)\rfloor$ and similarly, $r^{(s)}(\alpha)=\lfloor\log(1/\lvert I^{s}(\alpha^{T})\rvert)\rfloor$ for the stable scale of $\alpha$. Write $\alpha^{\ast}=(a_{1},a_{2},...,a_{n-1})$ if $\alpha=(a_{1},a_{2},...,a_{n})$ and for $r\in\mathbb{N}$ define the sets
@@
Note that by 2.2 for $\alpha\in\mathcal{C}_{u}(X,r)$ one has $e^{c_{1}}\lambda_{2}^{-1}\lambda_{2}^{\lvert\alpha\rvert}>\lvert I^{u}(\alpha^{\ast})\rvert>e^{-r}$ and from this follows that $\lvert\alpha\rvert<r/\log(\lambda_{2}^{-1})+\log(e^{c_{1}}\lambda_{2}^{-1})/\log(\lambda_{2}^{-1})$ and then
@@
where $\alpha_{1}=\log\lvert\mathcal{A}\rvert/\log(\lambda_{2}^{-1})>0$ and $\alpha_{2}=\log(e^{c_{1}}\lambda_{2}^{-1})\cdot\log\lvert\mathcal{A}\rvert/\log(\lambda_{2}^{-1})>0$ depends only on $\varphi$ and $\Lambda$. Note that the same inequality also holds for $N_{s}(X,r)$.
In the article [1] the authors proved the following lemma in the case of $X=\Lambda_{t}$, for completeness we give a proof here:
@@
By symmetry, it is suffices to show that the sequence $\{N_{u}(X,r)\}_{r\in\mathbb{N}}$ satisfies the conclusions of the lemma. By 2.1 and 2.2 we have for all $\alpha$, $\beta$, $\gamma$ finite words such that the concatenation $\alpha\beta\gamma$ is admissible
@@
$D_{u}(X)=\lim_{r\to\infty}\dfrac{\log N_{u}(X,r)}{r}=\inf\limits_{r\in\mathbb{N}}\dfrac{\log(|\mathcal{A}|^{c_{2}}\cdot N_{u}(X,r))}{r}$
@@
$D_{s}(X)=\lim_{r\to\infty}\dfrac{\log N_{s}(X,r)}{r}=\inf\limits_{r\in\mathbb{N}}\dfrac{\log(|\mathcal{A}|^{c_{2}}\cdot N_{s}(X,r))}{r}$
and that the numbers $D_{u}(X)$ and $D_{s}(X)$ are the limit capacities of $\pi^{u}(X)$ and $\pi^{s}(X)$ respectively.
By 2.2 we have for the constants $\tilde{C}=\log\lambda_{1}/\log\lambda_{2}>1$ and $C=e^{c_{1}\cdot(\tilde{C}+1)}>1$ and any $\alpha$ admissible that
@@
and for this, we conclude that for every $X\subset\Lambda$, compact and $\varphi$-invariant, $D_{s}(X)$ and $D_{u}(X)$ are comparable:

### v2

@@ Abstract.
Let $\varphi_{0}$ be a $C^{2}$-conservative diffeomorphism of a compact surface $S$ and let $\Lambda_{0}$ be a mixing horseshoe of $\varphi_{0}$. Given a smooth real function $f$ defined in $S$ and some diffeomorphism $\varphi$, close to $\varphi_{0}$, let $\mathcal{L}_{\varphi,f}$ be the Lagrange spectrum associated to the hyperbolic continuation $\Lambda(\varphi)$ of the horseshoe $\Lambda_{0}$ and $f$. We show that, for generic choices of $\varphi$ and $f$, if $L_{\varphi,f}$ is the map that gives the Hausdorff dimension of the set $\mathcal{L}_{\varphi,f}\cap(-\infty,t)$ for $t\in\mathbb{R}$, then there is an interval $I_{\varphi,f}$ which contains all the discontinuities of $L_{\varphi,f}$ with the end points of this interval the only possible limits of an infinite sequence of discontinuities.
@@ 1. Introduction
$\mathcal{L}:=\{\ell(\alpha):\alpha\in\mathbb{R}-\mathbb{Q}\ \text{and}\ \ell(\alpha)<\infty\}$
@@
Similarly, given a real quadratic form $q(x,y)=ax^{2}+bxy+cy^{2}$, let $\Delta(q)=b^{2}-4ac$ be its discriminant. We define the Markov spectrum as follows
$\mathcal{M}:=\left\{\frac{\sqrt{\Delta(q)}}{\inf\limits_{(x,y)\in\mathbb{Z}^{2}-\{(0,0)\}}|q(x,y)|}<\infty:q\ \text{is indefinite and}\ \Delta(q)>0\right\}.$
@@
  • Hall showed that $\mathcal{L}$ (and then $\mathcal{M}$) contains a half-line and Freiman determined the biggest half-line contained in the spectra, namely $c,+\infty)$ where
@@
For our purposes, it is worth to point out here that the Lagrange and Markov spectra have the following dynamical interpretation in terms of the continued fraction algorithm: Denote by $a_{0},a_{1},\dots]$ the continued fraction $a_{0}+\frac{1}{a_{1}+\frac{1}{\ddots}}$. Let $\Sigma=\mathbb{N}^{\mathbb{Z}}$ the space of bi-infinite sequences of positive integers, $\sigma:\Sigma\to\Sigma$ be the left-shift map $\sigma((a_{n})_{n\in\mathbb{Z}})=(a_{n+1})_{n\in\mathbb{Z}}$, and let $f:\Sigma\to\mathbb{R}$ be the function
@@
$\mathcal{L}=\left\{\limsup_{n\to\infty}f(\sigma^{n}(\underline{\theta}))<\infty:\underline{\theta}\in\Sigma\right\}\quad\textrm{and}\quad\mathcal{M}=\left\{\sup_{n\to\infty}f(\sigma^{n}(\underline{\theta}))<\infty:\underline{\theta}\in\Sigma\right\}.$
In the sequel, we consider the natural generalization of this dynamical version of the classical Lagrange and Markov spectra in the context of horseshoes of smooth diffeomorphisms of compact surfaces, where a horseshoe is a non-empty compact invariant hyperbolic set of saddle type which is transitive, locally maximal, and is not reduced to a periodic orbit.
@@ 1.2. Dynamical spectra
Let $\varphi:S\rightarrow S$ be a diffeomorphism of a $C^{\infty}$ compact surface $S$ with a mixing horseshoe $\Lambda$ and let $f:S\rightarrow\mathbb{R}$ be a differentiable function. Following the above characterization of the classical spectra, we define the maps $\ell_{\varphi,f}:\Lambda\rightarrow\mathbb{R}$ and $m_{\varphi,f}:\Lambda\rightarrow\mathbb{R}$ given by $\ell_{\varphi,f}(x)=\limsup\limits_{n\to\infty}f(\varphi^{n}(x))$ and $m_{\varphi,f}(x)=\sup\limits_{n\in\mathbb{Z}}f(\varphi^{n}(x))$ for $x\in\Lambda$ and call $\ell_{\varphi,f}(x)$ the Lagrange value of $x$ associated to $f$ and $\varphi$ and also $m_{\varphi,f}(x)$ the Markov value of $x$ associated to $f$ and $\varphi$. The sets
@@
are called Lagrange Spectrum of $(\varphi,f)$ and Markov Spectrum of $(\varphi,f)$. Here, we omitted the reference to the horseshoe $\Lambda$ because in our context it will always be determined by the diffeomorphism $\varphi.$
@@
In order to do that, we will explore the combinatorial nature of $\varphi|_{\Lambda}$ and its connection with the unstable and stable Cantor sets associated to $\Lambda$. More specifically, fix a Markov partition $\{R_{a}\}_{a\in\mathcal{A}}$ of $\Lambda$ with sufficiently small diameter consisting of rectangles $R_{a}\sim I_{a}^{u}\times I_{a}^{s}$ delimited by compact pieces $I_{a}^{s}$, $I_{a}^{u}$, of stable and unstable manifolds of certain points of $\Lambda$, see [16, Theorem 2, page 172]. The set $\mathcal{B}\subset\mathcal{A}^{2}$ of admissible transitions consists of pairs $(a,b)$ such that $\varphi(R_{a})\cap R_{b}\neq\emptyset$; so, we can define the transition matrix $B$ by
@@
Let $\Sigma_{\mathcal{A}}=\left\{\underline{a}=(a_{n})_{n\in\mathbb{Z}}:a_{n}\in\mathcal{A}\ \text{for all}\ n\in\mathbb{Z}\right\}$ and consider the homeomorphism of $\Sigma_{\mathcal{A}}$, the shift, $\sigma:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{A}}$ defined by $\sigma(\underline{a})_{n}=a_{n+1}$. Let $\Sigma_{\mathcal{B}}=\left\{\underline{a}\in\Sigma_{\mathcal{A}}:b_{a_{n}a_{n+1}}=1\right\}$, this set is closed and $\sigma$-invariant subspace of $\Sigma_{\mathcal{A}}$. Still denote by $\sigma$ the restriction of $\sigma$ to $\Sigma_{\mathcal{B}}$, the pair $(\Sigma_{\mathcal{B}},\sigma)$ is a subshift of finite type, see [6, Chapter 10]. The dynamics of $\varphi$ on $\Lambda$ is topologically conjugate to the sub-shift $\Sigma_{\mathcal{B}}$, namely, there is a homeomorphism $\Pi:\Lambda\to\Sigma_{\mathcal{B}}$ such that $\varphi\circ\Pi=\Pi\circ\sigma$.
As we will generally deal with sequences, we transfer the function $f$ from $\Lambda$ to a function (still denoted $f$) on $\Sigma_{\mathcal{B}}$. In this way, we set
@@
Recall that the stable and unstable manifolds of $\Lambda$ can be extended to locally invariant $C^{1+\alpha}$ foliations in a neighborhood of $\Lambda$ for some $\alpha>0$. Using these foliations it is possible to define projections $\pi^{u}_{a}:R_{a}\rightarrow I^{s}_{a}\times\{i^{u}_{a}\}$ and $\pi^{s}_{a}:R_{a}\rightarrow\{i^{s}_{a}\}\times I^{u}_{a}$ of the rectangles into the connected components $I^{s}_{a}\times\{i^{u}_{a}\}$ and $\{i^{s}_{a}\}\times I^{u}_{a}$ of the stable and unstable boundaries of $R_{a}$, where $i^{u}_{a}\in\partial I^{u}_{a}$ and $i^{s}_{a}\in\partial I^{s}_{a}$ are arbitrarily fixed. In this way, we have the unstable and stable Cantor sets
@@
Moreira’s theorem of [11] was generalized first in [1] in the context of conservative diffeomorphism with some horseshoe with Hausdorff dimension smaller than $1$, and later the condition on the dimension of the horseshoe was removed in [9]. More specifically, the authors proved that for typical choices of the conservative diffeomorphism and of the real function, the intersections of the corresponding dynamical Markov and Lagrange spectra with half-lines $(-\infty,t)$ have the same Hausdorff dimension, and this defines a continuous function of $t$ whose image is $0,\min\{1,D\}]$, where $D$ is the Hausdorff dimension of the horseshoe.
Our main theorem (cf. Theorem 1.1 below) is quite related to the result of the previous paragraph, but in our case, we consider typical diffeomorphism that are not necessarily conservative. Here, we will identify “the canonical interval” where $L_{\varphi,f}$ can have a discontinuity, and we will show that far away from the endpoints of this interval, one has finitely many points of discontinuity.
@@ 1.3. Statement of the main theorem
  • If $HD(\Lambda)<1$ then $L_{\varphi,f}$ has finitely many discontinuities in any closed subinterval $I\subset I_{\varphi,f}$ that does not contain $c_{\varphi,f}.$
  • If $HD(\Lambda)\geq 1$ then $L_{\varphi,f}$ has finitely many discontinuities in any closed subinterval $I\subset I_{\varphi,f}$ that does not contain neither $c_{\varphi,f}$ nor $\tilde{c}_{\varphi,f}.$
@@
If $HD(\Lambda_{0})<1$, then by choosing $\mathcal{U}$ small, given $\varphi\in\mathcal{U}^{*}$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and $\epsilon>0$ the function $L_{\varphi,f}$ has finitely many discontinuities in the interval $c_{\varphi,f}+\epsilon,\infty)$. Therefore, $c_{\varphi,f}$ is the only possible limit of an infinite sequence of discontinuities of $L_{\varphi,f}$.
@@
If $HD(\Lambda_{0})>1$, then by choosing $\mathcal{U}$ small, given $\varphi\in\mathcal{U}^{*}$, $f\in\mathcal{P}_{\varphi,\Lambda}$ and $\epsilon>0$ small, the function $L_{\varphi,f}$ has finitely many discontinuities in the interval $c_{\varphi,f}+\epsilon,\tilde{c}_{\varphi,f}-\epsilon].$ Therefore, $c_{\varphi,f}$ and $\tilde{c}_{\varphi,f}$ are the only possible limits of an infinite sequence of discontinuities of $L_{\varphi,f}$.
@@ 2.1. Stable and unstable dimensions
Note that if $x\in R(\theta;s)\cap\Lambda$, then the symbolic representation of $x$ is of the form $\Pi(x)=(\dots,a_{s_{1}}\dots a_{s-1};a_{s},a_{s+1}\dots a_{s_{2}}\dots)$, where the letter following to ; is in the 0 position of the sequence.
In our context of dynamically defined Cantor sets, we can relate the length of the unstable and stable intervals determined by an admissible word to its length as a word in the alphabet $\mathcal{A}$ via the bounded distortion property that lets us conclude that for some constant $c_{1}>0$
@@
[…]
@@
Note that by (2.2) for $\alpha\in\mathcal{C}_{u}(X,r)$ one has $e^{c_{1}}\lambda_{2}^{-1}\lambda_{2}^{\lvert\alpha\rvert}>\lvert I^{u}(\alpha^{\ast})\rvert>e^{-r}$ and it follows from this that $\lvert\alpha\rvert<r/\log(\lambda_{2}^{-1})+\log(e^{c_{1}}\lambda_{2}^{-1})/\log(\lambda_{2}^{-1})$ and then
@@
where $\alpha_{1}=\log\lvert\mathcal{A}\rvert/\log(\lambda_{2}^{-1})>0$ and $\alpha_{2}=\log(e^{c_{1}}\lambda_{2}^{-1})\cdot\log\lvert\mathcal{A}\rvert/\log(\lambda_{2}^{-1})>0$ depend only on $\varphi$ and $\Lambda$. This computations also give for $\alpha\in\mathcal{C}_{u}(X,r)$ that
(2.4) $\lvert\alpha\rvert<\alpha_{1}r+\alpha_{2}.$
Note that the same inequalities hold for $N_{s}(X,r)$ and words in $\mathcal{C}_{s}(X,r)$.
In the article [1] the authors proved the following lemma in the case of $X=\Lambda_{t}$, where $t\in\mathbb{R}$. For completeness, we give a proof here:
@@
By symmetry, it is sufficient to show that the sequence $\{N_{u}(X,r)\}_{r\in\mathbb{N}}$ satisfies the conclusions of the lemma. By (2.1) and (2.2) we have for all $\alpha$, $\beta$, $\gamma$ finite words such that the concatenation $\alpha\beta\gamma$ is admissible
@@
(2.5) $D_{u}(X)=\lim_{r\to\infty}\dfrac{\log N_{u}(X,r)}{r}=\inf\limits_{r\in\mathbb{N}}\dfrac{\log(|\mathcal{A}|^{c_{2}}\cdot N_{u}(X,r))}{r}$
@@
(2.6) $D_{s}(X)=\lim_{r\to\infty}\dfrac{\log N_{s}(X,r)}{r}=\inf\limits_{r\in\mathbb{N}}\dfrac{\log(|\mathcal{A}|^{c_{2}}\cdot N_{s}(X,r))}{r}.$
It can also be easily verified that the numbers $D_{u}(X)$ and $D_{s}(X)$ are the limit capacities of $\pi^{u}(X)$ and $\pi^{s}(X)$ respectively.
By (2.2) we have for the constants $\tilde{C}=\log\lambda_{1}/\log\lambda_{2}>1$ and $C=e^{c_{1}\cdot(\tilde{C}+1)}>1$ and any $\alpha$ admissible that
@@
and for this, we can conclude again that for every $X\subset\Lambda$, compact and $\varphi$-invariant, $D_{s}(X)$ and $D_{u}(X)$ are comparable: